# CityGaussianV2: Efficient and Geometrically Accurate Reconstruction for Large-Scale Scenes

### v1

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has revolutionized radiance field reconstruction, manifesting efficient and high-fidelity novel view synthesis. However, accurately representing surfaces, especially in large and complex scenarios, remains a significant challenge due to the unstructured nature of 3DGS. In this paper, we present CityGaussianV2, a novel approach for large-scale scene reconstruction that addresses critical challenges related to geometric accuracy and efficiency. Building on the favorable generalization capabilities of 2D Gaussian Splatting (2DGS), we address its convergence and scalability issues. Specifically, we implement a decomposed-gradient-based densification and depth regression technique to eliminate blurry artifacts and accelerate convergence. To scale up, we introduce an elongation filter that mitigates Gaussian count explosion caused by 2DGS degeneration. Furthermore, we optimize the CityGaussian pipeline for parallel training, achieving up to 10$\times$ compression, at least 25% savings in training time, and a 50% decrease in memory usage. We also established standard geometry benchmarks under large-scale scenes. Experimental results demonstrate that our method strikes a promising balance between visual quality, geometric accuracy, as well as storage and training costs. The project page is available at https://dekuliutesla.github.io/CityGaussianV2/.
@@ 1 Introduction
On the one hand, existing methods face significant challenges related to scalability and generalization ability. For example, SuGaR (Guédon & Lepetit, 2024) binds meshes with Gaussians for refinement. However, in large-scale and complex scenarios, meshes can become extensive and finely detailed, resulting in substantial memory demands and capacity limitations. GOF (Yu et al., 2024c) struggles with large, over-blurred Gaussians. These Gaussians not only severely deteriorate its rendering and geometry quality, but also generate a shell-like mesh that is non-trivial to remove, as validated in Fig. 7 and Fig. 8 in Appendix. While 2DGS (Huang et al., 2024) exhibits better generalization ability, as shown in Tab. 1, its convergence is hindered by the blurred Gaussians illustrated in part (b) of Fig. 1. Additionally, when scaling up through parallel training, it suffers from a Gaussian count explosion, as depicted in Fig. 3. Another challenge lies in the evaluation protocol: due to insufficient observations in boundary regions, geometry estimation becomes error-prone and unstable in these areas. As a result, the metrics can significantly fluctuate and underestimate actual performance (Xiong et al., 2024), making it difficult to objectively evaluate and compare algorithms.
On the other hand, achieving efficient parallel training and compression is critical to realizing geometrically accurate reconstruction of large-scale scenes. As shown in Tab. 2, the total number of Gaussians can increase to 19.3 million during parallel training, resulting in a storage requirement of 4.6 GB and a memory cost of 31.5 GB, while rendering speed drops below 25 FPS. Additionally, existing VastGaussian (Lin et al., 2024) costs nearly 3 hours for training, and CityGaussian Liu et al. (2024) consumes 4 hours to finish both training and compression. For reconstruction on low-end devices or under strict time constraints, these training costs and rendering speeds are unacceptable. Therefore, there is an urgent need for an economical parallel training and compression strategy.
@@ 3.1 Preliminary
3D Gaussian Splatting (Kerbl et al., 2023) represents 3D scene with a set of ellipsoids described by 3D Gaussian distribution, i.e. $\mathbf{G}_{\mathbf{N}}=\left\{G_{n}|n=1,...,N\right\}$. Each Gaussian contains learnable properties including central point $\bm{\mu}_{\bm{n}}\in\mathbb{R}^{3\times 1}$, covariance $\mathbf{\Sigma}_{\bm{n}}\in\mathbb{R}^{3\times 3}$, opacity $\sigma_{n}\in\left[0,1\right]$, spherical harmonics (SH) features $\bm{f}_{\bm{n}}\in\mathbb{R}^{3\times 16}$ for view-dependent rendering. The covariance matrix is further decomposed to scaling matrix $\mathbf{R}_{\bm{n}}$ and rotation matrix $\mathbf{S}_{\bm{n}}$, i.e. $\mathbf{\Sigma}_{\bm{n}}=\mathbf{R}_{\bm{n}}\mathbf{S}_{\bm{n}}{\mathbf{S}_{\bm{n}}}^{T}{\mathbf{R}_{\bm{n}}}^{T}$. For a certain pixel $p$, the color $\bm{c}_{p}$ is derived through alpha blending:
@@ 3.2 Optimization Mechanism
Figure 3: Illustration of the motivation and effectiveness of our Elongation Filter. We take the reconstruction of Rubble (Turki et al., 2022) scene as an example. On the left, we highlight the collection of Gaussian primitives with high gradient or extreme elongation. There is a significant overlap between two collections. By restricting densification of these sand-like points, we prevent out-of-memory (OOM) errors caused by an explosion in Gaussian count, enabling a steady count evolution analogous to CityGaussian (Liu et al., 2024) in parallel tuning, as depicted on the right.
@@ 3.3 Parallel Training Pipeline
As discussed in Sec. 1, the post-pruning and distillation of CityGaussian are time-consuming. And a significant memory overhead is introduced during parallel tuning. Considering these issues, we propose a novel pipeline, as shown in Fig. 4. To bypass the distillation step, we use an SH degree of 2 from the start, reducing the SH feature dimension from 48 to 27. This results in considerable memory and storage savings throughout the whole pipeline. To eliminate the need for post-pruning, we incorporate trimming during block-wise tuning. Specifically, we define the single-view contribution of each Gaussian following Fan et al. (2024):
@@ 4 Geometric evaluation protocols
To further deal with the challenge of boundary effect, an appropriate estimation of the crop volume is necessary. The core here is to check the visible frequency of each point and estimate a bound that can exclude rarely observed points. For efficiency, we take a workaround that formulates points as Gaussian primitives and checks their visibility using a well-optimized GS rasterizer. As illustrated in Fig. 5, we begin by initializing a 3DGS field with the ground-truth point cloud, then traverse all training views to rasterize and count visible frequency through the output visible mask. If the frequency of $j$-th point $\tau_{j}$ is below a predefined threshold, it will be excluded. The remaining points are then used to compute the alpha shape in the ground plane as well as the minimum and maximum heights. This process can be completed within 1 minute. Compared to the crop volume estimated on all points, ours reduces the error bar length of the F1 score from 0.1 to 0.003, enabling a stable, consistent, and reliable evaluation of the model’s actual performance.
@@ 5.1 Experimental Setup
Figure 6: Qualitative comparison of surface reconstruction quality. Here “Russian“ and “Modern“ denote the Russian Building and Modern Building scene of GauU-Scene, respectively. And “Aerial“ denotes aerial view of MatrixCity. The messy results of GOF are mainly attributed to the near-ground shell-like mesh. Due to the page limit, we leave further qualitative comparison with GOF and visualization at street view in the Appendix.
@@
Implementation Details. Our experiments are conducted on 8 A100 GPUs. We set the gradient scaling factor $\omega$ to 0.9 and the pruning ratio to 0.025. For depth distortion loss, we empirically find it harmful to performance, and thus set its weight to default value 0. The weight for $\mathcal{L}_{\mathrm{Depth}}$ is exponentially decayed from 0.5 to 0.0025 during both the pretraining and fine-tuning stages. $\mathcal{L}_{\mathrm{Normal}}$ is activated after 7,000 iterations in pretraining and from the beginning in the parallel tuning. Besides, we found that the original normal supervision was overly aggressive for complex scene reconstruction. Consequently, the weight for $\mathcal{L}_{\mathrm{Normal}}$ is reduced to 0.0125, one-fourth of its original value. We adhere to the default settings in CityGaussian (Liu et al., 2024) for the learning rate and densification schedule. To be specific, it applies lower learning rates during tuning compared to pertaining, and the street view is trained with a significantly lower learning rate and longer densification interval due to its extreme view sparsity (Zhou et al., 2024). Due to page limitations, additional details, including parameters for block partition and quantization, are provided in the Appendix.
@@ 5.2 Comparison with SOTA methods
Figure 7: Qualitative comparison of rendering quality. Here “Russian“ and “Modern“ denote the Russian Building and Modern Building scene of GauU-Scene, respectively. “Aerial“ denotes the aerial view of MatrixCity. Due to the page limit, we leave the street view in the Appendix.
Table 1: Comparison with SOTA reconstruction methods. “NaN“ means no results due to NaN error, and “OOM“ denotes out-of-memory error. “FAIL“ means the method fails to extract meaningful mesh due to poor convergence. Here P and R denotes precision and recall against ground-truth point cloud, respectively. More detailed rendering metrics are included in the Appendix.
@@
CityGS 24.75 0.457 0.371 0.407 27.46 0.362 0.637 0.462 22.98 0.283 0.689 0.401
Ours 24.51 0.576 0.450 0.501 27.23 0.441 0.752 0.556 22.19 0.376 0.759 0.503
In this section, we compare CityGaussianV2 with state-of-the-art (SOTA) methods both quantitatively and qualitatively. As shown in Tab. 1, NeRF-based methods are more prone to failure due to the NaN outputs of the MLP or poor convergence under sparse supervision in large-scale scenes. Besides, these methods generally take over 10 hours for training. In contrast, GS-based methods finish training within several hours, while demonstrating stronger performance and generalization abilities. For GauU-Scene, our model significantly outperforms existing geometry-specialized methods in rendering quality. As visually illustrated in Fig. 7, our method accurately reconstructs details such as crowded windows and woodlands. Geometrically, our model outperforms 2DGS by 0.01 F1 score. Besides, part (b) of Fig. 1 shows that even without parallel tuning, our proposed optimization strategy enables our model to achieve significantly better performance in rendering and geometry at both 7K and 30K iterations, while 2DGS struggles to efficiently optimize large and blurry surfels. This validates our superiority in convergence speed. Compared to CityGS, our method improves the F1 score by 0.1 while maintaining a minimal gap in PSNR. As validated in Fig. 6, the meshes produced by our method are smoother and more complete.
On the challenging MatrixCity dataset, we evaluate performance from both aerial and street views. For MatrixCity-Aerial, our method achieves the best surface quality among all algorithms, with the F1 score being twice that of 2DGS and outperforming CityGaussian by a significant margin. Furthermore, both SuGaR and GOF fail to complete training or extract meaningful meshes. In the street view, CityGS and geometry-specialized methods like 2DGS significantly underperform our method in geometry. As illustrated in Fig. 9 in the Appendix, our method provides qualitatively better reconstructions of road and building surfaces, with rendering quality comparable to CityGS.
Regarding training costs, as indicated in Tab. 2, the small version of CityGaussianV2 (ours-s) reduces training time by 25% and memory usage by over 50%, while delivering superior geometric performance and on-par rendering quality with CityGS. The tiny version (ours-t) can even halve the training time. These advantages make our method particularly suitable for scenarios with varying quality and immediacy requirements.
@@ Appendix A Additional Qualitative Comparison
This section provides additional qualitative comparisons. As illustrated in Fig. 8, the mesh produced by GOF is obscured by a near-ground shell, which obstructs rendering from the test view in Fig. 6 of the main paper and is challenging to remove. However, it does successfully capture the intricate structures of buildings and landscapes. For a more thorough comparison, we opt for the interior view in Fig. 6. Notably, our CityGaussianV2 showcases qualitatively better reconstructions with more geometry details and fewer outliers.
@@ Appendix B Additional Quantitative Results
Table 3: Detailed comparison with SOTA on rendering metrics. “NaN“ here means no results due to NaN error, and “OOM“ denotes no results due to out-of-memory error. “FAIL“ means fail to extract meaningful mesh due to poor convergence.
@@
Ours 0.765 24.51 0.215 0.501 0.857 27.23 0.169 0.531 0.784 22.19 0.382 0.524

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has revolutionized radiance field reconstruction, manifesting efficient and high-fidelity novel view synthesis. However, accurately representing surfaces, especially in large and complex scenarios, remains a significant challenge due to the unstructured nature of 3DGS. In this paper, we present CityGaussianV2, a novel approach for large-scale scene reconstruction that addresses critical challenges related to geometric accuracy and efficiency. Building on the favorable generalization capabilities of 2D Gaussian Splatting (2DGS), we address its convergence and scalability issues. Specifically, we implement a decomposed-gradient-based densification and depth regression technique to eliminate blurry artifacts and accelerate convergence. To scale up, we introduce an elongation filter that mitigates Gaussian count explosion caused by 2DGS degeneration. Furthermore, we optimize the CityGaussian pipeline for parallel training, achieving up to 10$\times$ compression, at least 25% savings in training time, and a 50% decrease in memory usage. We also established standard geometry benchmarks under large-scale scenes. Experimental results demonstrate that our method strikes a promising balance between visual quality, geometric accuracy, as well as storage and training costs. More live demos and official code implementation are available at our project page: https://dekuliutesla.github.io/CityGaussianV2/.
@@ 1 Introduction
On the one hand, existing methods face significant challenges related to scalability and generalization ability. For example, SuGaR (Guédon & Lepetit, 2024) binds meshes with Gaussians for refinement. However, it struggles to recover complex geometry details (Fig. 6) and can trigger out-of-memory errors when scaling up due to suboptimal implementation. GOF (Yu et al., 2024c) struggles with large, over-blurred Gaussians. These Gaussians obstruct the field of view and hinder valid supervision, leading to severe underfitting and shell-like mesh that is non-trivial to remove, as validated in Fig. 7 and Fig. 8 in Appendix. While 2DGS (Huang et al., 2024) exhibits better generalization ability, as shown in Tab. 1, its convergence is hindered by the blurred Gaussians illustrated in part (b) of Fig. 1. Additionally, when scaling up through parallel training, it suffers from a Gaussian count explosion, as depicted in Fig. 3. Another challenge lies in the evaluation protocol: due to insufficient observations in boundary regions, geometry estimation becomes error-prone and unstable in these areas. As a result, the metrics can significantly fluctuate and underestimate actual performance (Xiong et al., 2024), making it difficult to objectively evaluate and compare algorithms.
On the other hand, achieving efficient parallel training and compression is critical to realizing geometrically accurate reconstruction of large-scale scenes. The total number of Gaussians can increase to 19.3 million during parallel training, resulting in a storage requirement of 4.6 GB and a memory cost of 31.5 GB, while rendering speed drops below 25 FPS. Additionally, existing VastGaussian (Lin et al., 2024) costs nearly 3 hours for training, and CityGaussian Liu et al. (2024) consumes 4 hours to finish both training and compression. For reconstruction on low-end devices or under strict time constraints, these training costs and rendering speeds are unacceptable. Therefore, there is an urgent need for an economical parallel training and compression strategy.
@@ 3.1 Preliminary
3D Gaussian Splatting (Kerbl et al., 2023) represents 3D scene with a set of ellipsoids described by 3D Gaussian distribution, i.e. $\mathbf{G}_{\mathbf{N}}=\left\{G_{n}|n=1,...,N\right\}$. Each Gaussian contains learnable properties including central point $\bm{\mu}_{\bm{n}}\in\mathbb{R}^{3\times 1}$, covariance $\mathbf{\Sigma}_{\bm{n}}\in\mathbb{R}^{3\times 3}$, opacity $\sigma_{n}\in\left[0,1\right]$, spherical harmonics (SH) features $\bm{f}_{\bm{n}}\in\mathbb{R}^{3\times 16}$ for view-dependent rendering. The covariance matrix is further decomposed to scaling matrix $\mathbf{S}_{\bm{n}}$ and rotation matrix $\mathbf{R}_{\bm{n}}$, i.e. $\mathbf{\Sigma}_{\bm{n}}=\mathbf{R}_{\bm{n}}\mathbf{S}_{\bm{n}}{\mathbf{S}_{\bm{n}}}^{T}{\mathbf{R}_{\bm{n}}}^{T}$. For a certain pixel $p$, the color $\bm{c}_{p}$ is derived through alpha blending:
@@ 3.2 Optimization Mechanism
Figure 3: Illustration of the motivation and effectiveness of our Elongation Filter. We take the tuning of one block of Rubble (Turki et al., 2022) scene as an example. On the left, we highlight the collection of Gaussian primitives with high gradient or extreme elongation. There is a significant overlap between two collections. By restricting densification of these sand-like points, we prevent out-of-memory (OOM) errors caused by an explosion in Gaussian count, enabling a steady count evolution analogous to CityGaussian (Liu et al., 2024) in parallel tuning, as depicted on the right.
@@ 3.3 Parallel Training Pipeline
As discussed in Sec. 1, the post-pruning and distillation of CityGaussian leads to time and memory overhead. To resolve these issues, we propose a novel pipeline, as shown in Fig. 4. To bypass the distillation step, we use an SH degree of 2 from the start, reducing the SH feature dimension from 48 to 27. This results in considerable memory and storage savings throughout the whole pipeline. To eliminate the need for post-pruning, we incorporate trimming during block-wise tuning. Specifically, we define the single-view contribution of each Gaussian following Fan et al. (2024):
@@ 4 Geometric evaluation protocols
To further deal with the challenge of boundary effect, an appropriate estimation of the crop volume is necessary. The core here is to check the visible frequency of each point and estimate a bound that can exclude rarely observed points. For efficiency, we take a workaround that formulates points as Gaussian primitives and checks their visibility using a well-optimized GS rasterizer. As illustrated in Fig. 5, we begin by initializing a 3DGS field with the ground-truth point cloud, then traverse all training views to rasterize and count visible frequency through the output visible mask. If the frequency of $j$-th point $\tau_{j}$ is below a predefined threshold, it will be excluded. Then we calculate the minimum and maximum height of the remaining points, and project them to the ground plane with ground-truth transformation matrix for alpha shape estimation. Given a scene covering $1.47km^{2}$ with 958 training views and 31.4 million ground-truth points, this process can be completed within 1 minute if rendered in 1080p on a 40G A100. Compared to the crop volume estimated on all points, ours reduces the error bar length of the F1 score from 0.1 to 0.003, enabling a stable, consistent, and reliable evaluation of the model’s actual performance.
@@ 5.1 Experimental Setup
Figure 6: Qualitative comparison of surface reconstruction quality. Here “Russian“ and “Modern“ denote the Russian Building and Modern Building scene of GauU-Scene, respectively. And “Aerial“ denotes aerial view of MatrixCity. The messy results of GOF are mainly attributed to the near-ground shell-like mesh. More visualizations about GOF and street view scene are in the Appendix.
@@
Implementation Details. All experiments included in this paper are conducted on 8 A100 GPUs. We set the gradient scaling factor $\omega$ to 0.9 and the pruning ratio to 0.025. For depth distortion loss, we empirically find it harmful to performance, and thus set its weight to default value 0. The weight for $\mathcal{L}_{\mathrm{Depth}}$ is exponentially decayed from 0.5 to 0.0025 during both the pretraining and fine-tuning stages. $\mathcal{L}_{\mathrm{Normal}}$ is activated after 7,000 iterations in pretraining and from the beginning in the parallel tuning. Besides, we found that the original normal supervision was overly aggressive for complex scene reconstruction. Consequently, the weight for $\mathcal{L}_{\mathrm{Normal}}$ is reduced to 0.0125, one-fourth of its original value. We adhere to the default settings in CityGaussian (Liu et al., 2024) for the learning rate and densification schedule. Due to page limitations, detailed parameters for block partition and quantization are provided in the Appendix.
@@ 5.2 Comparison with SOTA methods
Figure 7: Qualitative comparison of rendering quality. Here “Russian“ and “Modern“ denote the Russian Building and Modern Building scene of GauU-Scene, respectively. “Aerial“ denotes the aerial view of MatrixCity. The result on street view is included in Fig. 9 of the Appendix.
Table 1: Comparison with SOTA reconstruction methods. “NaN“ means no results due to NaN error. “FAIL“ means the method fails to extract meaningful mesh due to poor convergence. Here P and R denotes precision and recall against ground-truth point cloud, respectively.
@@
CityGS 24.75 0.522 0.405 0.453 27.46 0.362 0.637 0.462 22.98 0.283 0.689 0.401
Ours 24.51 0.576 0.450 0.501 27.23 0.441 0.752 0.556 22.24 0.376 0.759 0.503
In this section, we compare CityGaussianV2 with state-of-the-art (SOTA) methods both quantitatively and qualitatively. Tab. 1 report results with no compression. As shown, NeRF-based methods are more prone to failure due to the NaN outputs of the MLP or poor convergence under sparse supervision in large-scale scenes. Besides, these methods generally take over 10 hours for training. In contrast, GS-based methods finish training within several hours, while demonstrating stronger performance and generalization abilities. For GauU-Scene, our model significantly outperforms existing geometry-specialized methods in rendering quality. As visually illustrated in Fig. 7, our method accurately reconstructs details such as crowded windows and woodlands. Geometrically, our model outperforms 2DGS by 0.01 F1 score. Besides, part (b) of Fig. 1 shows that even without parallel tuning, our proposed optimization strategy enables our model to achieve significantly better performance in rendering and geometry at both 7K and 30K iterations, while 2DGS struggles to efficiently optimize large and blurry surfels. This validates our superiority in convergence speed. Compared to CityGS, though 0.24 PSNR is sacrificed, our method gains around 11% F1-score improvement. As validated in Fig. 6, our meshes are smoother and more complete.
On the challenging MatrixCity dataset, we evaluate performance from both aerial and street views. For MatrixCity-Aerial, our method achieves the best surface quality among all algorithms, with the F1 score being twice that of 2DGS and outperforming CityGaussian by a significant margin. Furthermore, GOF fails to complete training or extract meaningful meshes. In the street view, CityGS and geometry-specialized methods like 2DGS significantly underperform our method in geometry. As illustrated in Fig. 9 in the Appendix, our method provides qualitatively better reconstructions of road and building surfaces, with rendering quality comparable to CityGS.
Regarding training costs, as indicated in Tab. 2, the small version of CityGaussianV2 (ours-s) reduces training time by 25% and memory usage by over 50%, while delivering superior geometric performance and on-par rendering quality with CityGS. The tiny version (ours-t) can even halve the training time. These advantages make our method particularly suitable for scenarios with varying quality and immediacy requirements. Results on other scenes are included in Tab. 4 of Appendix.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by the National Key R&D Program of China (No. 2022ZD0116500), the National Natural Science Foundation of China (No. U21B2042, No. 62320106010), and in part by the 2035 Innovation Program of CAS.
@@ Appendix A Additional Qualitative Comparison
Figure 11: Bad cases visualization. Due to occlusion and lack of observation, some road surfaces and building facades are not well reconstructed. TSDF-based fusion also struggles to recover some thin structures like spires.
This section provides additional qualitative comparisons. As illustrated in Fig. 8, the mesh produced by GOF is obscured by a near-ground shell, which obstructs rendering from the test view in Fig. 6 of the main paper and is challenging to remove. However, it does successfully capture the intricate structures of buildings and landscapes. Fig. 8 provides a more thorough comparison. Notably, our CityGaussianV2 showcases qualitatively better reconstructions with more geometry details and fewer outliers.
@@ Appendix B Additional Quantitative Results
Table 3: Detailed comparison with SOTA on rendering metrics. “NaN“ here means no results due to NaN error. “FAIL“ means fail to extract meaningful mesh due to poor convergence.
@@
Ours 0.765 24.51 0.215 0.501 0.857 27.23 0.169 0.531 0.788 22.24 0.347 0.524
Table 4: Detailed comparison among SOTA among parallel training methods. 2DGS* here means applying CityGS’s training strategy to 2DGS without our proposed optimization mechanism. And “OOM“ means one or more sub-blocks fail to finish training due to the out-of-memory error. The best result for specific metrics under each scene is highlighted in bold.
[…]
2DGS* OOM OOM OOM OOM OOM OOM OOM
Residence CityGS 23.17 0.453 8.05 235 0.44 31.5 66.7
Ours 23.46 0.465 8.07 181 0.44 14.2 45.5
2DGS* OOM OOM OOM OOM OOM OOM OOM
Russia CityGS 24.19 0.455 7.00 209 0.38 27.4 55.2
Ours 23.89 0.537 6.97 177 0.38 15.0 33.3
2DGS* OOM OOM OOM OOM OOM OOM OOM
Modern CityGS 26.22 0.462 7.90 215 0.43 29.2 57.1
Ours 25.53 0.489 7.90 185 0.42 16.1 34.5
2DGS* OOM OOM OOM OOM OOM OOM OOM
Aerial CityGS 27.23 0.459 10.3 217 0.56 25.7 38.6
Ours 26.70 0.492 10.4 181 0.56 14.8 27.0
2DGS* 22.24 0.371 9.20 170 2.17 15.5 28.6
Street CityGS 21.12 0.398 7.63 163 0.42 11.9 50.0
Ours 22.09 0.499 7.59 149 0.42 10.8 31.3
@@
In addition to the full experimental results in Tab. 1 and Tab. 3, we provide a comparison of parallel training methods with compressed and aligned Gaussian counts in Tab. 4. For reference, we include results where 2DGS is directly paired with CityGS’s parallel training strategy. However, as shown, this approach encounters OOM errors in most scenes due to excessive memory demands caused by redundant Gaussians and the Gaussian count explosion issue illustrated in Fig. 3. Compared to CityGS, our method (the small version, i.e. ours-s in Tab. 2) achieves superior geometric accuracy while significantly reducing training time and memory usage. Under extreme compression (e.g., 75% on Residence) or in street-view scenes, our method also delivers significantly better rendering quality. These results not only highlight the necessity of our proposed optimization strategy but also demonstrate our method’s clear advantages over CityGS.